# Minimum Selective Subset on Unit Disk Graphs and Circle Graphs

### v2

@@ Abstract
We show that the MSS problem is log-APX-hard on general graphs, even when $c=2$. As a consequence, the problem does not admit a polynomial-time approximation scheme (PTAS) unless $\textup{{P}}=\textup{{NP}}$. On the positive side, we present a PTAS for unit disk graphs that does not require a geometric representation and applies for arbitrary $c$. We further prove that MSS remains NP-complete in unit disk graphs for arbitrary $c$. In addition, we show that the MSS problem is APX-hard on circle graphs, even when $c=2$.
@@ 1 Introduction
Many computational tools have been developed for supervised learning methods on a labeled training set $T$ embedded in a metric space $(X,d)$. Each data point $t\in T$ is associated with a label (also called a character or color), chosen from a set $C=\{1,2,\dots,c\}$. The objective is to extract a smallest possible subset $S\subseteq T$ such that every point in $T$ either belongs to $S$ or has at least one nearest neighbor (with respect to the metric $d$) within $S$ that shares the same character. This optimization problem, called the Minimum Consistent Subset (MCS), was originally formulated by Hart [Hart] in 1968, which has received thousands of citations, highlighting its significant impact in the field. However, the paper [Hart] did not establish any complexity results or algorithms.
@@
Later in 1991, Wilfong [Wilfong] defined two problems MCS and MSS together and proved that the MCS and MSS problems are NP-complete in $\mathbb{R}^{2}$ for $c\geq 3$ and $c\geq 2$, respectively. It also proposed a polynomial-time algorithm when there is only one red point and all other points are blue in $\mathbb{R}^{2}$. Later, in 2018, it was proved that MCS remains NP-complete when $c=2$ in $\mathbb{R}^{2}$ [Khodamoradi]. Recently, Banerjee et al. [BBC] showed that MCS is W[2]-hard (for arbitrary $c$) and MSS is W[1] (for $c=2$), both parameterized by the solution size. Various algorithms, including those for many restricted inputs for the MCS problem in $\mathbb{R}^{2}$ have been proposed [BBC, Ahmad, Chitnis22], highlighting its significance in machine learning and computational geometry. The only algorithm for the MSS problem is a PTAS, which was established when $c=2$ [BBC].
@@
The Minimum Selective Subset (MSS) problem plays a crucial role in optimizing data selection by identifying the smallest subset which preserves essential information. It can be viewed both as a clustering and a proximity problem. This is particularly useful in applications such as fingerprint recognition, character recognition, and pattern recognition, where it helps reduce redundancy and improve decision-making in classification and feature selection tasks. MSS was introduced because the standard MCS methods, like Hart’s [Hart], could not guarantee that the resulting subset was the smallest possible size. MSS applies a stronger consistency condition on the subset to create a mathematical framework that is amenable to an exact minimization procedure, thereby fulfilling the optimization goal that MCS heuristics failed to satisfy. So far, we have discussed MCS and MSS problems along with their published results in $\mathbb{R}^{2}$. We now turn to these problems in the context of graph algorithms.
@@
Banerjee et al. [BBC] proved that MCS is W[2]-hard [book] when parameterized by the solution size, even with only two colors on general graphs. Dey et al. [DeyMN21, DeyMN23] provided polynomial-time algorithms for MCS on some simple graph classes including paths, spiders, caterpillars, combs, and trees (for trees, $c=2$). XP, NP-complete, and FPT (when $c$ is a parameter) results on trees, can be found in [Arimura23, aritra]. The MCS problem is also NP-complete on interval graphs [aritra] and APX-hard on circle graphs [2024arXiv240514493M]. Variants, such as the Minimum Consistent Spanning Subset (MCSS) and the Minimum Strict Consistent Subset (MSCS) of MCS, have been studied on trees [banik2024minimum, biniaz2024minimum, manna2024minimumstrictconsistentsubset]. However, the algorithmic results for MSS have not been extensively studied to date. Banerjee et al. [BBC] only showed that MSS is NP-complete on general graphs. Very recently, the MSS problem has been studied in various settings, including $\mathcal{O}(\log n)$-approximation algorithms for general graphs, NP-complete results for planar graphs, and linear-time algorithms for trees and unit interval graphs [manna2025minimumselectivesubsetgraph] (published in CCCG 2025).
@@
Our Contributions. The MSS problem admits an $\mathcal{O}(\log n)$-approximation on general graphs, which raises the question of whether better approximations exist. We show in Section 3 that MSS is log-APX-hard even when $c=2$. Hence, the problem is also APX-hard and does not admit a PTAS on general graphs. This leads to the natural question of whether some graph classes allow a PTAS. To date, none are known. We answer this by proving in Section 5 that MSS admits a PTAS on unit disk graphs for arbitrary $c$, without requiring a geometric representation. Unit disk graphs are also fundamental in wireless networks, robotics, and computational geometry, where efficient approximation algorithms are highly relevant [GOLDIN2009234]. Before presenting our PTAS, we establish in Section 4 that MSS remains NP-complete on unit disk graphs when $c$ is arbitrary. We also investigate whether MSS is APX-hard in other graph classes. In Section 6, we prove that MSS is APX-hard on circle graphs even when $c=2$. Circle graphs, which model intersecting chords, have applications in VLSI design, scheduling, and bioinformatics [article, sherwani2012algorithms]. All proofs of results marked with $(^{\ast})$ can be found in the Appendix.
@@ 2 Preliminaries
Let $G=(V(G),E(G))$ be a graph, where $V(G)$ is the vertex set and $E(G)$ is the edge set. For any $U\subseteq V(G)$, $G[U]$ denotes the subgraph of $G$ induced on $U$, and $\lvert U\rvert$ is the cardinality of $U$. We denote $[n]$ as the set of integers $\{1,\ldots,n\}$. We use an arbitrary vertex color function $C:V(G)\rightarrow[c]$, which assigns each vertex exactly one color from the set $[c]$. For a subset of vertices $U\subseteq V(G)$, let $C(U)$ represent the set of colors of the vertices in $U$, formally defined as $C(U)=\{C(u)\mid u\in U\}$. The shortest path distance (i.e., hop-distance) between two vertices $u$ and $v$ in $G$ is denoted by $\mathrm{d}(u,v)$. Distance between $v\in V(G)$ and the set $U\subseteq V(G)$ is given by $\mathrm{d}(v,U)=\min_{u\in U}\mathrm{d}(v,u)$. Similarly, the distance between two subgraphs $G_{1}$ and $G_{2}$ in $G$ is defined as $\mathrm{d}(G_{1},G_{2})=\min\{\mathrm{d}(v_{1},v_{2})\mid v_{1}\in V(G_{1}),v_{2}\in V(G_{2})\}$. The set of nearest neighbors of $v$ in the set $U$ is denoted as $\mathrm{\hat{N}}(v,U)$, formally defined as $\mathrm{\hat{N}}(v,U)=\{u\in U\mid\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. Therefore, if $v\in U$, then $\mathrm{\hat{N}}(v,U)=\{v\}$. The set of vertices in $U$ adjacent to $v$ is given by $\mathrm{N}(v,U)=\{u\in U\mid(u,v)\in E(G)\}.$ We also define $\mathrm{N}[v,U]=\{v\}\cup\mathrm{N}(v,U).$ For any two subsets $U_{1},U_{2}\subseteq V(G)$, we define $\mathrm{N}(U_{1},U_{2})=\bigcup_{v\in U_{1}}\mathrm{N}(v,U_{2}),$ and $\mathrm{N}[U_{1},U_{2}]=\bigcup_{v\in U_{1}}\mathrm{N}[v,U_{2}].$ Most symbols and notations follow standard conventions from [diestel2012graph]. Suppose $G=(V(G),E(G))$ is a given simple connected undirected graph where $\bigcup_{i=1}^{c}V_{i}=V(G)$ and $V_{i}\cap V_{j}=\emptyset$ for $i\neq j$ and each vertex in $V_{i}$ is assigned color $i$. A Minimum Consistent Subset (MCS) is a subset $S\subseteq V(G)$ of minimum cardinality such that for every vertex $v\in V(G)$, if $v\in V_{i}$, then $\mathrm{\hat{N}}(v,S)\cap V_{i}$ is non-empty.
@@
We establish a reduction from the Minimum Dominating Set (MDS) problem to the MSS problem. In the Minimum Dominating Set problem, the input is a graph $G$ together with an integer $s$. The task is to decide whether there exists a subset $D\subseteq V(G)$ of size at most $s$ such that every vertex $u\in V(G)$ satisfies $\mathrm{N}[u,V(G)]\cap D\neq\emptyset$. It is well known that the Minimum Set Cover problem is log-APX-hard (see [10.5555/1965254] for complexity class definitions), and moreover, it is NP-hard to approximate it within a factor of $\delta\cdot\log n$ for some positive constant $\delta$ [raz97]. Since there exists an $L$-reduction from the Minimum Set Cover problem to the Minimum Dominating Set problem, the Minimum Dominating Set problem is also log-APX-hard.
@@ 4 NP-completeness of MSS on Unit Disk Graphs
Clark et al. [Clark1990] showed that Minimum Dominating Set (MDS) is NP-complete in UDG. We reduce from an instance $U$ of the UDG with $\lvert V(U)\rvert=n$ to an instance $U^{\prime}$ as follows:
@@ 5.3 Finding $S^{min}(E_{i,1}^{j})$ on Unit Disk Graphs
Now, we apply the method described in [independent] to find a Maximum Independent Set in $F_{i,j}$ for unit disk graphs.
@@ 6 APX-hardness of MSS on Circle Graphs
The vertex set of a circle graph is a set of chords of a given circle, and if two chords intersect, the corresponding vertices share an edge. We obtain a “gap-preserving” reduction from the MAX-3SAT(8) problem to a circle graph using the Minimum Dominating Set (MDS) problem on circle graphs. The MAX-3SAT(8) problem is as follows [10.5555/1965254]:
We are given a set of $n$ variables $\mathcal{X}=\{x_{1},\dots,x_{n}\}$ and $m$ clauses $\mathcal{C}=\{c_{1},\dots,c_{m}\}$ such that each clause has at most $3$ literals and each variable occurs in at most 8 clauses. The objective is to find a truth-assignment of the variables in $\mathcal{X}$ that maximizes the number of clauses in $\mathcal{C}$ satisfied. MAX-3SAT(8) is APX-hard, and the MDS problem on circle graphs is also APX-hard [circle].
Consider an instance $\phi$ of MAX-3SAT(8). Let $\text{SAT}(\phi)$ represent the maximum fraction of clauses in $\phi$ that can be satisfied. For a given graph $G$, let $\gamma(G)$ denote the cardinality of its Minimum Dominating Set. The paper [circle] reduces an instance of MAX-3SAT(8) to a circle graph $G$ such that $\lvert V(G)\rvert=m+56n+4$ and states the following theorem:
@@ Appendix 0.D Proof of the Lemma 3
It is straightforward to see that MSS belongs to NP, since a given solution can be verified in polynomial time. By Lemma 3, we established a reduction from the NP-complete MDS problem in UDG [Clark1990]. Hence, MSS is NP-complete on unit disk graphs.∎
@@ Appendix 0.Q Proof of the Lemma 12
As established in [GrahamKnuthPatashnik1998], for $0\leq\epsilon<\frac{1}{2}$, the following inequality holds

### v3

# On Computing Minimum Selective Subsets in Unit Disk and Circle Graphs

@@ Abstract
We show that the MSS problem is log-APX-hard on general graphs, even when $c=2$. As a consequence, the problem does not admit a polynomial-time approximation scheme (PTAS) unless $\textup{{P}}=\textup{{NP}}$. On the positive side, we present a PTAS for unit disk graphs that does not require a geometric representation and applies for arbitrary $c$. We further prove that MSS remains NP-complete in unit disk graphs for arbitrary $c$. In addition, we show that the MSS problem is APX-hard on circle graphs, even when $c=2$  [^1] This work appeared in CALDAM 2026 [21].
@@ 1 Introduction
Many computational tools have been developed for supervised learning methods on a labeled training set $T$ embedded in a metric space $(X,d)$. Each data point $t\in T$ is associated with a label (also called a character or color), chosen from a set $C=\{1,2,\dots,c\}$. The objective is to extract a smallest possible subset $S\subseteq T$ such that every point in $T$ either belongs to $S$ or has at least one nearest neighbor (with respect to the metric $d$) within $S$ that shares the same character. This optimization problem, called the Minimum Consistent Subset (MCS), was originally formulated by Hart [16] in 1968, which has received thousands of citations, highlighting its significant impact in the field. However, the paper [16] did not establish any complexity results or algorithms.
@@
Later in 1991, Wilfong [27] defined two problems MCS and MSS together and proved that the MCS and MSS problems are NP-complete in $\mathbb{R}^{2}$ for $c\geq 3$ and $c\geq 2$, respectively. It also proposed a polynomial-time algorithm when there is only one red point and all other points are blue in $\mathbb{R}^{2}$. Later, in 2018, it was proved that MCS remains NP-complete when $c=2$ in $\mathbb{R}^{2}$ [17]. Recently, Banerjee et al. [2] showed that MCS is W[2]-hard (for arbitrary $c$) and MSS is W[1] (for $c=2$), both parameterized by the solution size. Various algorithms, including those for many restricted inputs for the MCS problem in $\mathbb{R}^{2}$ have been proposed [2, 5, 7], highlighting its significance in machine learning and computational geometry. The only algorithm for the MSS problem is a PTAS, which was established when $c=2$ [2].
@@
The Minimum Selective Subset (MSS) problem plays a crucial role in optimizing data selection by identifying the smallest subset which preserves essential information. It can be viewed both as a clustering and a proximity problem. This is particularly useful in applications such as fingerprint recognition, character recognition, and pattern recognition, where it helps reduce redundancy and improve decision-making in classification and feature selection tasks. MSS was introduced because the standard MCS methods, like Hart’s [16], could not guarantee that the resulting subset was the smallest possible size. MSS applies a stronger consistency condition on the subset to create a mathematical framework that is amenable to an exact minimization procedure, thereby fulfilling the optimization goal that MCS heuristics failed to satisfy. So far, we have discussed MCS and MSS problems along with their published results in $\mathbb{R}^{2}$. We now turn to these problems in the context of graph algorithms.
@@
Banerjee et al. [2] proved that MCS is W[2]-hard [9] when parameterized by the solution size, even with only two colors on general graphs. Dey et al. [11, 12] provided polynomial-time algorithms for MCS on some simple graph classes including paths, spiders, caterpillars, combs, and trees (for trees, $c=2$). XP, NP-complete, and FPT (when $c$ is a parameter) results on trees, can be found in [1, 4]. The MCS problem is also NP-complete on interval graphs [4] and APX-hard on circle graphs [18]. Variants, such as the Minimum Consistent Spanning Subset (MCSS) and the Minimum Strict Consistent Subset (MSCS) of MCS, have been studied on trees [3, 6, 19]. However, the algorithmic results for MSS have not been extensively studied to date. Banerjee et al. [2] only showed that MSS is NP-complete on general graphs. Very recently, the MSS problem has been studied in various settings, including $\mathcal{O}(\log n)$-approximation algorithms for general graphs, NP-complete results for planar graphs, and linear-time algorithms for trees and unit interval graphs [20] (published in CCCG 2025).
@@
Our Contributions. The MSS problem admits an $\mathcal{O}(\log n)$-approximation on general graphs, which raises the question of whether better approximations exist. We show in Section 3 that MSS is log-APX-hard even when $c=2$. Hence, the problem is also APX-hard and does not admit a PTAS on general graphs. This leads to the natural question of whether some graph classes allow a PTAS. To date, none are known. We answer this by proving in Section 5 that MSS admits a PTAS on unit disk graphs for arbitrary $c$, without requiring a geometric representation. Unit disk graphs are also fundamental in wireless networks, robotics, and computational geometry, where efficient approximation algorithms are highly relevant [14]. Before presenting our PTAS, we establish in Section 4 that MSS remains NP-complete on unit disk graphs when $c$ is arbitrary. We also investigate whether MSS is APX-hard in other graph classes. In Section 6, we prove that MSS is APX-hard on circle graphs even when $c=2$. Circle graphs, which model intersecting chords, have applications in VLSI design, scheduling, and bioinformatics [26, 24]. All proofs of results marked with $(^{\ast})$ can be found in the Appendix.
@@ 2 Preliminaries
Let $G=(V(G),E(G))$ be a graph, where $V(G)$ is the vertex set and $E(G)$ is the edge set. For any $U\subseteq V(G)$, $G[U]$ denotes the subgraph of $G$ induced on $U$, and $\lvert U\rvert$ is the cardinality of $U$. We denote $[n]$ as the set of integers $\{1,\ldots,n\}$. We use an arbitrary vertex color function $C:V(G)\rightarrow[c]$, which assigns each vertex exactly one color from the set $[c]$. For a subset of vertices $U\subseteq V(G)$, let $C(U)$ represent the set of colors of the vertices in $U$, formally defined as $C(U)=\{C(u)\mid u\in U\}$. The shortest path distance (i.e., hop-distance) between two vertices $u$ and $v$ in $G$ is denoted by $\mathrm{d}(u,v)$. Distance between $v\in V(G)$ and the set $U\subseteq V(G)$ is given by $\mathrm{d}(v,U)=\min_{u\in U}\mathrm{d}(v,u)$. Similarly, the distance between two subgraphs $G_{1}$ and $G_{2}$ in $G$ is defined as $\mathrm{d}(G_{1},G_{2})=\min\{\mathrm{d}(v_{1},v_{2})\mid v_{1}\in V(G_{1}),v_{2}\in V(G_{2})\}$. The set of nearest neighbors of $v$ in the set $U$ is denoted as $\mathrm{\hat{N}}(v,U)$, formally defined as $\mathrm{\hat{N}}(v,U)=\{u\in U\mid\mathrm{d}(v,u)=\mathrm{d}(v,U)\}$. Therefore, if $v\in U$, then $\mathrm{\hat{N}}(v,U)=\{v\}$. The set of vertices in $U$ adjacent to $v$ is given by $\mathrm{N}(v,U)=\{u\in U\mid(u,v)\in E(G)\}.$ We also define $\mathrm{N}[v,U]=\{v\}\cup\mathrm{N}(v,U).$ For any two subsets $U_{1},U_{2}\subseteq V(G)$, we define $\mathrm{N}(U_{1},U_{2})=\bigcup_{v\in U_{1}}\mathrm{N}(v,U_{2}),$ and $\mathrm{N}[U_{1},U_{2}]=\bigcup_{v\in U_{1}}\mathrm{N}[v,U_{2}].$ Most symbols and notations follow standard conventions from [13]. Suppose $G=(V(G),E(G))$ is a given simple connected undirected graph where $\bigcup_{i=1}^{c}V_{i}=V(G)$ and $V_{i}\cap V_{j}=\emptyset$ for $i\neq j$ and each vertex in $V_{i}$ is assigned color $i$. A Minimum Consistent Subset (MCS) is a subset $S\subseteq V(G)$ of minimum cardinality such that for every vertex $v\in V(G)$, if $v\in V_{i}$, then $\mathrm{\hat{N}}(v,S)\cap V_{i}$ is non-empty.
@@
We establish a reduction from the Minimum Dominating Set (MDS) problem to the MSS problem. In the Minimum Dominating Set problem, the input is a graph $G$ together with an integer $s$. The task is to decide whether there exists a subset $D\subseteq V(G)$ of size at most $s$ such that every vertex $u\in V(G)$ satisfies $\mathrm{N}[u,V(G)]\cap D\neq\emptyset$. It is well known that the Minimum Set Cover problem is log-APX-hard (see [25] for complexity class definitions), and moreover, it is NP-hard to approximate it within a factor of $\delta\cdot\log n$ for some positive constant $\delta$ [23]. Since there exists an $L$-reduction from the Minimum Set Cover problem to the Minimum Dominating Set problem, the Minimum Dominating Set problem is also log-APX-hard.
@@ 4 NP-completeness of MSS on Unit Disk Graphs
Clark et al. [8] showed that Minimum Dominating Set (MDS) is NP-complete in UDG. We reduce from an instance $U$ of the UDG with $\lvert V(U)\rvert=n$ to an instance $U^{\prime}$ as follows:
@@ 5.3 Finding $S^{min}(E_{i,1}^{j})$ on Unit Disk Graphs
Now, we apply the method described in [22] to find a Maximum Independent Set in $F_{i,j}$ for unit disk graphs.
@@ 6 APX-hardness of MSS on Circle Graphs
The vertex set of a circle graph is a set of chords of a given circle, and if two chords intersect, the corresponding vertices share an edge. We obtain a “gap-preserving” reduction from the MAX-3SAT(8) problem to a circle graph using the Minimum Dominating Set (MDS) problem on circle graphs. The MAX-3SAT(8) problem is as follows [25]:
We are given a set of $n$ variables $\mathcal{X}=\{x_{1},\dots,x_{n}\}$ and $m$ clauses $\mathcal{C}=\{c_{1},\dots,c_{m}\}$ such that each clause has at most $3$ literals and each variable occurs in at most 8 clauses. The objective is to find a truth-assignment of the variables in $\mathcal{X}$ that maximizes the number of clauses in $\mathcal{C}$ satisfied. MAX-3SAT(8) is APX-hard, and the MDS problem on circle graphs is also APX-hard [10].
Consider an instance $\phi$ of MAX-3SAT(8). Let $\text{SAT}(\phi)$ represent the maximum fraction of clauses in $\phi$ that can be satisfied. For a given graph $G$, let $\gamma(G)$ denote the cardinality of its Minimum Dominating Set. The paper [10] reduces an instance of MAX-3SAT(8) to a circle graph $G$ such that $\lvert V(G)\rvert=m+56n+4$ and states the following theorem:
@@ Appendix 0.D Proof of the Lemma 3
It is straightforward to see that MSS belongs to NP, since a given solution can be verified in polynomial time. By Lemma 3, we established a reduction from the NP-complete MDS problem in UDG [8]. Hence, MSS is NP-complete on unit disk graphs.∎
@@ Appendix 0.Q Proof of the Lemma 12
As established in [15], for $0\leq\epsilon<\frac{1}{2}$, the following inequality holds